# Simulating Molecular Single Vibronic Level Fluorescence Spectra with ab initio Hagedorn Wavepacket Dynamics

### v4

@@ 4 Results and Discussion
Figure 1: SVL fluorescence spectra of anthracene from initial vibrational levels $12^{j}$ and $\overline{11}^{j}$ ($j=1,2$) computed from Hagedorn wavepacket dynamics in the adiabatic harmonic model; a scaling factor of 0.97 was applied to the wavenumbers of the computed spectra (red dashes); the experimental reference (black solid line) is taken from ref 3.
@@
Anthracene belongs to the $D_{2h}$ point group. Within the Condon approximation, only vibrational levels belonging to the $\mathrm{a_{g}}$ representation can be prepared in the excited $\mathrm{{}^{1}B_{2u}}$ state from the ground vibrational level of the ground $\mathrm{{}^{1}A_{g}}$ state to serve as the initial states in SVL experiments.3, 14 However, through Herzberg–Teller intensity borrowing, the preparation of SVL initial states with $\mathrm{b_{1g}}$ vibrational symmetry is also possible. As examples, we show below the SVL spectra with initial vibrational excitations in modes $12$ (of $\mathrm{a_{g}}$ symmetry) and $\overline{11}$ (of $\mathrm{b_{1g}}$ symmetry).
@@
Figure 1 compares the spectra of anthracene evaluated with the Hagedorn approach in the adiabatic harmonic model to the experimental emission spectra from the $12^{1}$, $12^{2}$, $\overline{11}^{1}$, and $\overline{11}^{2}$ levels.3 For these SVL spectra, the Franck–Condon selection rules for emissions provide that transitions from $\mathrm{a_{g}}$ levels ($12^{1}$, $12^{2}$, and $\overline{11}^{2}$) are only allowed to $\mathrm{a_{g}}$ levels, and that transitions from the $\mathrm{b_{1g}}$ level ($\overline{11}^{1}$) are only allowed to $\mathrm{b_{1g}}$ levels. The total quantum numbers in $\mathrm{b_{1g}}$ modes must thus retain the same parity (odd or even) after the transition.
@@
In fig 1, the computed and experimental SVL fluorescence spectra are shown with respect to the wavenumber differences $\tilde{\nu}_{\text{exc}}-\tilde{\nu}_{\text{em}}$ between the initial excitation light and the emitted light. An empirical scaling factor was applied to the wavenumber differences of the simulated spectra since the density functional theory (DFT) calculations systematically overestimate the vibrational frequencies of the ground electronic state.80, 81, 14 By optimizing the alignment of peak positions between the simulated and experimental spectra, we set the scaling factor to 0.97, which is close to the vibrational scaling factors reported for the PBE0 functional with similar basis sets.82
@@
The simulated spectra from both singly excited levels (first row in fig 1) agree well with the experiment and are consistent with the results obtained by Tapavicza using a generating function approach.14 The overall structure of the $12^{1}$ spectrum is reproduced, including the progression of mode 12 ($312^{1}_{j},j=0,1,2,$ 3) and the transitions involving a mode change to another $\mathrm{a_{g}}$ mode (e.g., $12^{1}_{0}6^{0}_{1}$) or combinations with other $\mathrm{a_{g}}$ modes (e.g., $12^{1}_{1}6^{0}_{1}$ and $12^{1}_{2}6^{0}_{1}$). While the intensity of the $12^{1}_{0}$ peak is overestimated, the computed intensities of certain combination bands (for example, $12^{1}_{1}7^{0}_{1}$, $12^{1}_{1}6^{0}_{1}$ and $12^{1}_{2}6^{0}_{1}$) are significantly lower than experimentally observed. From the $\overline{11}^{1}$ level, the $\overline{11}^{1}_{0}$ transition is negligible in the experimental spectrum and entirely absent in the computed spectrum, as expected from symmetry considerations. Whereas the harmonic Hagedorn dynamics reproduces the cluster of combination bands between 1500 and 2500 $\text{cm}^{-1}$, the peaks of $\overline{11}^{1}_{0}\overline{10}^{0}_{1}$, $\overline{11}^{1}_{0}\overline{10}_{1}^{0}12_{1}^{0}$ , and $\overline{11}^{1}_{0}\overline{10}_{1}^{0}6_{1}^{0}$ transitions are severely underestimated or missing in the computed spectrum. The harmonic model allows clear peak assignments in the computed spectra. For example, while the $\overline{11}^{1}_{0}\overline{10}_{1}^{0}6_{1}^{0}$ and $\overline{11}^{1}_{1}4_{1}^{0}$ peaks are overlapping in the experiment, the peak present at 1954 $\text{cm}^{-1}$ (2015 $\text{cm}^{-1}$ before wavenumber scaling) in the computed $\overline{11}^{1}$ spectrum can be unambiguously assigned to the $\overline{11}^{1}_{1}4_{1}^{0}$ instead of $\overline{11}^{1}_{0}\overline{10}_{1}^{0}6_{1}^{0}$ transition based on the ab initio vibrational frequencies.
@@
In contrast to the alternative method from ref 14, the Hagedorn approach makes it possible to compute SVL spectra from multiply excited levels and using the same trajectory as the ground-level emission. The ab initio Hagedorn results for levels $12^{2}$ and $\overline{11}^{2}$ (second row in fig 1) agree reasonably well with the experiments. The computed $12^{2}$ spectrum correctly captures the experimentally observed decrease (compared to the $12^{1}$ spectrum) in the intensities of the $12^{j}_{0}$, $12^{j}_{j}$, and $12^{j}_{j}6^{0}_{1}$ peaks.
@@
As demonstrated in ref 42, the Hagedorn wavepacket dynamics is exact in harmonic potentials and can treat mode distortion and mode mixing (Duschinsky rotation) exactly. Therefore, the differences between experimental and simulated spectra must be due to the inaccuracies of the global harmonic model from electronic structure calculations, neglect of Herzberg–Teller contributions, significant anharmonicity of the true potential energy surface, or possibly (but unlikely) experimental error. To better understand the discrepancies, we show below the effects of different global harmonic approximations and the influence of Duschinsky coupling on the computed SVL spectra of anthracene. The importance of anharmonicity is also examined by comparing spectra from vertical, adiabatic, and local harmonic approximations.
Figure 2: Effects of displacement, mode distortion, and Duschinsky rotation on the simulated (red dashes) $12^{2}$ SVL fluorescence spectra of anthracene; a scaling factor of 0.97 was applied to the wavenumbers of the computed spectra. All spectra are compared to experiment (ref 3, black solid line).
@@
The displacement of the equilibrium position of the ground-state surface from the excited-state equilibrium produces more interesting, nontrivial spectral features (second row of fig 2). In the vertically and horizontally displaced harmonic oscillator model, the wavepacket was propagated with a potential centered around the ground-state equilibrium geometry but whose Hessian is identical to the excited-state Hessian (i.e., $\kappa_{g}=\kappa_{e}$ in $V_{g}$). Several experimental features (e.g., $12^{2}_{0}$, $12^{2}_{1}$, $12^{2}_{2}$, $12^{2}_{1}6^{0}_{1}$, and $12^{2}_{2}6^{0}_{1}$) are now reproduced, but some peaks (e.g., $12^{2}_{3}$, $12^{2}_{4}$, and $12^{2}_{3}6^{0}_{1}$) are underestimated or not resolved in the computed spectrum. The description of the $12^{2}_{3}$, $12^{2}_{4}$, and $12^{2}_{3}6^{0}_{1}$ peaks is improved in the third row by including mode distortion in the ground state (though the intensity of the $12^{2}_{2}6^{0}_{1}$ peak becomes worse). Mode distortion, which allows frequency changes between the ground- and excited-state surfaces, was included by defining the ground-state potential using only the diagonal components of the computed ground-state Hessian at the ground-state equillibrium geometry. As the off-diagonal Hessian elements were set to zero, the mode mixing effects were completely neglected in the computed spectrum shown in the third row.
@@
Figure 3: Effects of an artificially enhanced Duschinsky coupling between modes $\overline{11}$ and $\overline{10}$ on the computed $\overline{11}^{1}$ and $\overline{11}^{2}$ SVL fluorescence spectra of anthracene; a scaling factor of 0.97 was applied to the wavenumbers of the computed spectra (red dashes); the experimental reference (black solid line) is taken from ref 3.
@@
Figure 4: $12^{2}$ SVL fluorescence spectra of anthracene evaluated with the vertical (left, green dashes) and local harmonic (right, blue dashes) approaches. The spectra are compared to the adiabatic harmonic (first row, red solid line) and the experimental (ref 3, second row, black solid line) spectra; a scaling factor of 0.97 was applied to all computed spectra.
@@
The need for empirical wavenumber scaling in our results (see the Supporting Information for corresponding spectra without wavenumber scaling) indicates that anharmonicity may also be important. To test the adequacy of the global harmonic model, we first computed spectra with the vertical harmonic model where the ground-state surface is expanded around the excited-state equilibrium geometry (Franck–Condon point) instead of the ground-state minimum energy point. If the potential energy surface were compeletly harmonic, the two different harmonic approximations should give identical results. Figure 4 shows that the $12^{2}$ SVL spectrum obtained from the vertical harmonic model has notable differences in peak intensities and positions compared to the adiabatic harmonic spectrum (top left). Although the overall structure remains similar, the adiabatic harmonic spectrum (bottom left, fig 1) provided better agreement with the experiment than the vertical harmonic one (bottom left, fig 4).
@@
To confirm the impact of anharmonicity in anthracene, we also performed the much more expensive on-the-fly ab initio local harmonic calculations61. The local harmonic approximation produced very similar results to the global adiabatic harmonic approach (top right, fig 4) and empirical wavenumber scaling is still required to achieve good agreement with experimental wavenumbers (bottom right). This suggests that the (adiabatic) harmonic model suffices for the evaluation of SVL spectra of anthracene and that the empirical wavenumber scaling is primarily compensating for systematic bias (e.g., due to electronic correlation) in the chosen DFT method. The sensitivity of the computed spectra to the different harmonic models in figs 2 to 4 reflects how vibronic spectra may reveal detailed information about molecular potential energy surfaces in experiments.
@@
Our method can treat not only multiple excitations in a single mode, but also initial states where several vibrational modes are simultaneously excited to any vibrational level. Here, we have restricted ourselves to the experimentally accessible vibronic levels ($12^{1}6^{1}$, $12^{1}\overline{5}^{1}$, and $12^{1}5^{1}$), which were reported respectively at $1380+385\,$cm-1, $1409+385\,$cm-1, and $1420+385\,$cm-1 in the experimental fluorescence excitation spectrum of anthracene, albeit with very weak intensities (2–3% of the signal at origin).3 Despite the current lack of experimental data on the SVL fluorescence from these levels for a direct comparison, the Hagedorn approach can serve as a tool for predicting complex vibronic spectra.
@@
Figure 5 shows the SVL emission spectra from the levels $12^{1}6^{1}$, $12^{1}\overline{5}^{1}$, and $12^{1}5^{1}$ predicted with the adiabatic harmonic model. These spectra are broadly similar to the simulated $12^{1}$ spectrum, but the excitation in an additional mode leads to different intensity patterns. Compared to the $12^{1}6^{1}$ spectrum, the predicted spectrum from the $12^{1}5^{1}$ level shows weaker intensities in the first three peaks corresponding to a progression in mode 12. In constrast to $12^{1}6^{1}$ and $12^{1}5^{1}$ spectra, in the $12^{1}\overline{5}^{1}$ spectrum the three peaks of the mode 12 progression under 100010001000\,1000cm-1 (i.e., $12^{1}_{j}\overline{5}^{1}_{0}$ with $j=0,1,2$) disappear, because transitions to $\mathrm{a_{g}}$ levels (e.g., the ground level and $12_{j}$ levels) in the ground state are forbidden for initial levels of $\mathrm{b_{1g}}$ symmetry. The progression in mode 12 is only allowed when coupled with the symmetry-allowed $\overline{5}^{1}_{1}$ transition (e.g., $12_{0}^{1}\overline{5}_{1}^{1}$, $12_{1}^{1}\overline{5}_{1}^{1}$, and $12_{2}^{1}\overline{5}_{1}^{1}$). Otherwise, the features above 250025002500\,2500cm-1 (mainly combination bands with other $\mathrm{a_{g}}$ modes, analogous to the ${>}1000\,$cm-1 peaks in the $12^{1}$ spectrum) are similar in the SVL spectra from $12^{1}\overline{5}^{1}$ and $12^{1}5^{1}$ levels.

### v5

@@ 4 Results and Discussion
Figure 1: SVL fluorescence spectra of anthracene from initial vibrational levels $12^{j}$ and $\overline{11}^{j}$ ($j=1,2$) computed from Hagedorn wavepacket dynamics in the adiabatic harmonic model; a scaling factor of 0.97 was applied to the wavenumbers of the computed spectra (red dashed line); the experimental reference (black solid line) is taken from ref 3.
@@
Anthracene belongs to the $D_{2h}$ point group. Within the Condon approximation, only vibrational levels belonging to the $\mathrm{a_{g}}$ representation can be prepared in the excited $\mathrm{{}^{1}B_{2u}}$ state from the ground vibrational level of the ground $\mathrm{{}^{1}A_{g}}$ state to serve as the initial states in SVL experiments.3, 14 However, through Herzberg–Teller intensity borrowing, the preparation of SVL initial states with $\mathrm{b_{1g}}$ vibrational symmetry is also possible. Out of the 66 vibrational modes, twelve have $\mathrm{a_{g}}$ symmetry and eleven have $\mathrm{b_{1g}}$ symmetry (denoted by a bar in the mode label). As examples, we show below the SVL spectra with initial vibrational excitations in the $\mathrm{a_{g}}$ mode 12 and $\mathrm{b_{1g}}$ mode $\overline{11}$.
@@
Figure 1 compares the spectra of anthracene evaluated with the Hagedorn approach in the adiabatic harmonic model to the experimental emission spectra from the $^{1}, 1221$$^{2}$, $\overline{11}^{1}$, and $\overline{11}^{2}$ levels.3 For these SVL spectra, the Franck–Condon selection rules for emissions provide that transitions from $\mathrm{a_{g}}$ levels ($^{1}, 1221$$^{2}$, and $\overline{11}^{2}$) are only allowed to $\mathrm{a_{g}}$ levels, and that transitions from the $\mathrm{b_{1g}}$ level ($\overline{11}^{1}$) are only allowed to $\mathrm{b_{1g}}$ levels. The total quantum numbers in $\mathrm{b_{1g}}$ modes must thus retain the same parity (odd or even) after the transition.
@@
In fig 1, the computed and experimental SVL fluorescence spectra are shown with respect to the wavenumber differences $\tilde{\nu}_{\text{exc}}-\tilde{\nu}_{\text{em}}$ between the initial excitation light and the emitted light. An empirical scaling factor was applied to the wavenumber differences of the simulated spectra since the density functional theory (DFT) calculations systematically overestimate the vibrational frequencies of the ground electronic state.86, 87, 14 By optimizing the alignment of peak positions between the simulated and experimental spectra, we set the scaling factor to 0.97, which is close to the vibrational scaling factors reported for the PBE0 functional with similar basis sets.88
@@
The simulated spectra from both singly excited levels (first row in fig 1) agree well with the experiment and are consistent with the results obtained by Tapavicza using a generating function approach (see fig S2 in the Supporting Information). Both Tapavicza’s approach and our Hagedorn method are exact in harmonic models and are thus equivalent for singly excited levels. The overall structure of the $12^{1}$ spectrum is reproduced, including the progression of mode 12 ($12^{1}_{j},j=0,1,2,3$) and the transitions involving a mode change to another $\mathrm{a_{g}}$ mode (e.g., $12^{1}_{0}6^{0}_{1}$) or combinations with other $\mathrm{a_{g}}$ modes (e.g., $12^{1}_{1}6^{0}_{1}$ and $12^{1}_{2}6^{0}_{1}$). While the intensity of the $12^{1}_{0}$ peak is overestimated, the computed intensities of certain combination bands (for example, $12^{1}_{1}7^{0}_{1}$, $12^{1}_{1}6^{0}_{1}$ and $12^{1}_{2}6^{0}_{1}$) are significantly lower than experimentally observed. From the $\overline{11}^{1}$ level, the $\overline{11}^{1}_{0}$ transition is negligible in the experimental spectrum and entirely absent in the computed spectrum, as expected from symmetry considerations. Whereas the harmonic Hagedorn dynamics reproduces the cluster of combination bands between 1500 and 2500 $\text{cm}^{-1}$, the peaks of $\overline{11}^{1}_{0}\overline{10}^{0}_{1}$, $\overline{11}^{1}_{0}\overline{10}_{1}^{0}12_{1}^{0}$ , and $\overline{11}^{1}_{0}\overline{10}_{1}^{0}6_{1}^{0}$ transitions are severely underestimated or missing in the computed spectrum. The harmonic model allows clear peak assignments in the computed spectra. For example, while the $\overline{11}^{1}_{0}\overline{10}_{1}^{0}6_{1}^{0}$ and $\overline{11}^{1}_{1}4_{1}^{0}$ peaks are overlapping in the experiment, the peak present at 1954 $\text{cm}^{-1}$ (2015 $\text{cm}^{-1}$ before wavenumber scaling) in the computed $\overline{11}^{1}$ spectrum can be unambiguously assigned to the $\overline{11}^{1}_{1}4_{1}^{0}$ instead of $\overline{11}^{1}_{0}\overline{10}_{1}^{0}6_{1}^{0}$ transition based on the ab initio vibrational frequencies.
@@
In contrast to the alternative method from ref 14, the Hagedorn approach makes it possible to compute SVL spectra from multiply excited levels and using the same trajectory as that already needed for the ground-level emission. The ab initio Hagedorn results for levels $12^{2}$ and $\overline{11}^{2}$ (second row in fig 1) agree reasonably well with the experiments. The computed $12^{2}$ spectrum correctly captures the experimentally observed decrease (compared to the $12^{1}$ spectrum) in the intensities of the $12^{j}_{0}$, $12^{j}_{j}$, and $12^{j}_{j}6^{0}_{1}$ peaks.
@@
As demonstrated in ref 43, the Hagedorn wavepacket dynamics is exact in harmonic potentials and can treat mode distortion (frequency changes) and mode mixing (Duschinsky rotation) exactly. Therefore, the differences between experimental and simulated spectra must be due to the inaccuracies of the global harmonic model from electronic structure calculations, neglect of Herzberg–Teller contributions (particularly for $\mathrm{b_{1g}}$ levels), significant anharmonicity of the true potential energy surface, or possibly (but unlikely) experimental error. To better understand the discrepancies, we show below the effects of different global harmonic approximations and the influence of Duschinsky coupling on the computed SVL spectra of anthracene. The importance of anharmonicity is also examined by comparing spectra from vertical, adiabatic, and local harmonic approximations.
Figure 2: Effects of displacement, mode distortion, and mode mixing on the simulated (red dashed line) $12^{2}$ SVL fluorescence spectra of anthracene; a scaling factor of 0.97 was applied to the wavenumbers of the computed spectra. All spectra are compared to experiment (ref 3, black solid line).
@@
The displacement of the equilibrium position of the ground-state surface from the excited-state equilibrium produces more interesting, nontrivial spectral features (second row of fig 2). In the vertically and horizontally displaced harmonic oscillator model, the wavepacket was propagated with a potential centered around the ground-state equilibrium geometry but whose Hessian is identical to the excited-state Hessian (i.e., $\kappa_{g}=\kappa_{e}$ in $V_{g}$). Several experimental features (e.g., $^{2}_{0}, 1212$$^{2}_{1}$, $^{2}_{2}, 1212$$^{2}_{1}6^{0}_{1}$, and $12^{2}_{2}6^{0}_{1}$) are now reproduced, but some peaks (e.g., $12^{2}_{3}$, $12^{2}_{4}$, and $12^{2}_{3}6^{0}_{1}$) are underestimated or not resolved in the computed spectrum. The description of the $12^{2}_{3}$, $12^{2}_{4}$, and $12^{2}_{3}6^{0}_{1}$ peaks is improved in the third row by including mode distortion in the ground state (though the intensity of the $12^{2}_{2}6^{0}_{1}$ peak becomes worse). Mode distortion, which allows frequency changes between the ground- and excited-state surfaces, was included by defining the ground-state potential using only the diagonal components of the computed ground-state Hessian at the ground-state equillibrium geometry. As the off-diagonal Hessian elements were set to zero, the mode mixing effects were completely neglected in the computed spectrum shown in the third row.
@@
Figure 3: Effects of an artificially enhanced Duschinsky coupling between modes $\overline{11}$ and $\overline{10}$ on the computed $\overline{11}^{1}$ and $\overline{11}^{2}$ SVL fluorescence spectra of anthracene; a scaling factor of 0.97 was applied to the wavenumbers of the computed spectra (red dashed line); the experimental reference (black solid line) is taken from ref 3.
@@
Figure 4: $12^{2}$ SVL fluorescence spectra of anthracene evaluated with the vertical (left, green dashed line) and local harmonic (right, blue dashed line) approaches. The spectra are compared to the adiabatic harmonic (first row, red solid line) and the experimental (ref 3, second row, black solid line) spectra; a scaling factor of 0.97 was applied to all computed spectra.
@@
The need for empirical wavenumber scaling in our results (see the Supporting Information for corresponding spectra without wavenumber scaling) indicates that anharmonicity may also be important. To test the adequacy of the global harmonic model, we first computed spectra with the vertical harmonic model where the ground-state surface is expanded around the excited-state equilibrium geometry (Franck–Condon point) instead of the ground-state minimum energy point. If the potential energy surface were completely harmonic, the two different harmonic approximations should give identical results. Figure 4 shows that the $12^{2}$ SVL spectrum obtained from the vertical harmonic model has notable differences in peak intensities and positions compared to the adiabatic harmonic spectrum (top left). Although the overall structure remains similar, the adiabatic harmonic spectrum (bottom left, fig 1) provided better agreement with the experiment than the vertical harmonic one (bottom left, fig 4).
To confirm the impact of anharmonicity in anthracene, we also performed the much more expensive on-the-fly ab initio local harmonic calculations.62 Here, we expand the potential to second order based on DFT evaluations of gradients and Hessians at the current position, rather than using a state-independent global harmonic potential. The center of the wavepacket then follows a fully anharmonic classical trajectory, while the evolution of the width (matrices $Q_{t}$ and $P_{t}$) is determined by the local Hessian at each time step. The local harmonic approximation produced very similar results to the global adiabatic harmonic approach (top right, fig 4). Figure 5 compares the position evolution in mode 12 for the three approximations. The adiabatic and local harmonic trajectories are nearly identical with small differences in the amplitude of motion, whereas the vertical harmonic model gives a significantly different trajectory in the amplitude and frequency of oscillation. This suggests that the (adiabatic) harmonic model suffices for the evaluation of SVL spectra of anthracene.
Figure 5: Evolution of the position along mode 12 ($q_{\text{mode 12}}$) of the center of the wavepacket in the on-the-fly local harmonic dynamics (blue solid line), compared to the vertical (left panel, green dashed line) and adiabatic (right panel, red dashed line) harmonic dynamics.
@@
However, empirical wavenumber scaling is still required for the local harmonic results to achieve good agreement with experimental wavenumbers (bottom right, fig 4). Although the use of scaling factors are typically attributed to anharmonicity, our results suggest they may be primarily compensating for systematic bias (e.g., due to electronic correlation) in the chosen DFT method. While outside the scope of the present work focused on the dynamics method, further studies using anharmonic frequency corrections (e.g., with perturbation theory89, 90) could help separate electronic structure errors from genuine anharmonic contributions. The sensitivity of the computed spectra to the different harmonic models in figs 2 to 4 reflects how vibronic spectra may reveal detailed information about molecular potential energy surfaces in experiments.
We also note that the additional cost of computing overlaps for different initial states is negligible compared to propagation in the local harmonic case. As a rough comparison, in the global harmonic model, propagating the parameters of the Gaussian for 100,000 steps took approximately 12 CPU minutes (single-core) in total, while computing 25,000 overlaps for a doubly excited level required around 72 CPU minutes on the same computer. Whereas constructing the global harmonic potential for anthracene required only a single electronic structure evaluation (approximately 90 CPU minutes), on-the-fly local harmonic dynamics required performing such a calculation at each time step.
@@
Since the overlap expressions for two Hagedorn functions are general for any non-negative multi-index $K$, our method can use the same trajectory to treat not only multiple excitations in a single mode, but also initial states where several vibrational modes are simultaneously excited to any vibrational level (although the cost of computing overlaps increases as more modes are excited to higher levels). Here, we have restricted ourselves to the experimentally accessible vibronic levels ($12^{1}6^{1}$, $12^{1}\overline{5}^{1}$, and $12^{1}5^{1}$), which were reported respectively at 1380+3851380+385\,cm-1, 1409+3851409+385\,cm-1, and 1420+3851420+385\,cm-1 in the experimental fluorescence excitation spectrum of anthracene, albeit with very weak intensities (2–3% of the signal at origin).3 Despite the current lack of experimental data on the SVL fluorescence from these levels for a direct comparison, the Hagedorn approach can serve as a tool for predicting complex vibronic spectra.
@@
Figure 6 shows the SVL emission spectra from the levels $12^{1}6^{1}$, $12^{1}\overline{5}^{1}$, and $12^{1}5^{1}$ predicted with the adiabatic harmonic model. These spectra are broadly similar to the simulated $12^{1}$ spectrum, but the excitation in an additional mode leads to different intensity patterns. Compared to the $12^{1}6^{1}$ spectrum, the predicted spectrum from the $12^{1}5^{1}$ level shows weaker intensities in the first three peaks corresponding to a progression in mode 12. In constrast to $12^{1}6^{1}$ and $12^{1}5^{1}$ spectra, in the $12^{1}\overline{5}^{1}$ spectrum the three peaks of the mode 12 progression under 10001000\,cm-1 (i.e., $12^{1}_{j}\overline{5}^{1}_{0}$ with $j=0,1,2$) disappear, because transitions to $\mathrm{a_{g}}$ levels (e.g., the ground level and $12_{j}$ levels) in the ground state are forbidden for initial levels of $\mathrm{b_{1g}}$ symmetry. The progression in mode 12 is only allowed when coupled with the symmetry-allowed $\overline{5}^{1}_{1}$ transition (e.g., $12_{0}^{1}\overline{5}_{1}^{1}$, $12_{1}^{1}\overline{5}_{1}^{1}$, and $12_{2}^{1}\overline{5}_{1}^{1}$). Otherwise, the features above 25002500\,cm-1 (mainly combination bands with other $\mathrm{a_{g}}$ modes, analogous to the >1000{>}1000\,cm-1 peaks in the $12^{1}$ spectrum) are similar in the SVL spectra from $12^{1}\overline{5}^{1}$ and $12^{1}5^{1}$ levels.